Far detector installation commissioning operation
This presentation is the property of its rightful owner.
Sponsored Links
1 / 43

Far Detector Installation, Commissioning, & Operation PowerPoint PPT Presentation


  • 89 Views
  • Uploaded on
  • Presentation posted in: General

Far Detector Installation, Commissioning, & Operation. Jeff Nelson South Carolina Meeting March, 2003. Outline. FD Assembly Planes, staffing, & status Shield update FD Commissioning Front ends, LI, DAQ, DCS, Magnet Documentation update Data quality & operations update

Download Presentation

Far Detector Installation, Commissioning, & Operation

An Image/Link below is provided (as is) to download presentation

Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author.While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server.


- - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - -

Presentation Transcript


Far detector installation commissioning operation

Far Detector Installation, Commissioning, & Operation

Jeff Nelson

South Carolina Meeting

March, 2003


Outline

Outline

  • FD Assembly

    • Planes, staffing, & status

    • Shield update

  • FD Commissioning

    • Front ends, LI, DAQ, DCS, Magnet

    • Documentation update

    • Data quality & operations update

    • Electronic echoes (Roy Lee’s talk)

  • Planning for shifts & operations

  • CDMS 2 - progress worth noting


Far detector installation commissioning operation

J. Meier

1/30/03


Plane assembly

Plane Assembly

  • 426 planes installed

    • 86% completed, 4.75 kton

    • Averaging 5.5 PPW for SM2(was 6.4 in SM1) due to

    • Directly related to FTEs worked

      • Vacation, departures, just-in-time delivery mode for much of SM2

  • Some numbers

    • Steel at Soudan: 90%

    • Modules at Soudan: 94%

    • Over 1100 loads down shaft

    • 60 planes to go

    • 85th week of installation


Schedule projection

Schedule Projection

  • We are planning our schedule based on the plane construction completing June 2nd

    • Current projection is May 28th

    • Future based on some additional departures…


Staffing future projections

Staffing & Future Projections

  • Crew in flux

    • 7 departures since last summer (of the 22 original builders/movers)

    • We have hired 3 replacements

    • 3 more temps will start in March

  • Expect to lose more to the mines

    • 50 more hires at just one mine this month

    • Many of the new workers are from construction trades

      • May leave when the road restrictions lift this spring

      • Some may leave sooner as construction starts

    • Could be as many as 5 to 7 more departures

  • There will also be a substantial demobilization by materials handling crew after the last steel/scintillator is underground

  • This leaves a +2/-1 week uncertainty in projections


More schedule

More Schedule

  • Roughly 2 months to complete the coil, shield, & final detector shake down

  • After the last large objects are removed in very early June we will go to a days-only schedule

    • Maybe 9-hour days summer and maybe 8 hours days fall/winter

    • Significantly reduced crew staffing next summer


Scinticicles

Scinticicles

  • Scintillator shipped in early February had some problems

    • Some of the skins delaminated

    • Some strips show difference in readout

    • These were the first modules shipped in very cold weather

  • 19 modules from that shipment mapped again

    • 4 modules with problems (15 strips)

      • 2 connector, 4 manifold, 7 fibers unglued (2 more unclassified)

      • Ones with strange initial maps appear more vulnerable

    • Some may be installed before we noticed

    • Have to compare LI to maps to know

  • These are the first major differences in scintillator response observed at Soudan

    • Now sample mapping and transmission testing CIT modules from Dec and late Feb shipments


Some strips in the worst module

Some strips in the worst module


Commissioning

Commissioning

  • 417 planes reading out

    • Last commissioning on Friday

    • Mostly routine at this point


Shield update

Shield Update

  • Since December we commissioned the revised Section 1 shield

    • Shorter cables and repositioned modules

  • Added lower walls to sections 1 & 2 (next page)

  • Future work on hold until funding situation is clarified

  • When/if OK’d

    • Can install section 3 now

    • Can reconfigure section 2 as double layer

  • See Caius’s talk for the justification


Shield lower wall installation

Shield Lower Wall Installation


Fd daq rack hardware

FD DAQ/Rack Hardware

  • Crate commissioning

    • Crate 15 (the last crate)

      • All the marginal bits ended up here

      • Tass finally debugged

    • Some damage in things commissioned but not used for a few months

      • Swapped with spares (aka ND DAQ)

      • Not sure the effect

      • Significant modifications to the crate procedures (pull cards for extensive work; sign off before power)

  • Branches (PVICs)

    • Vendor has a hardware fix to the node/length problem

    • Swap through spares for upgrades after an initial test of a four node branch at Soudan

    • Months and a number of expert visits before we convert back to 4 branches


Daq software at soudan areas of emphasis

DAQ Software at SoudanAreas of Emphasis

  • ROP configuration overrides (Tim&Tass)

    • Override of dynode threshold settings for a scan

    • Override of sparsification mode to allow unsparsified running

    • Override of VARC trigger to allow untriggered running

    • Masks so that high rate special runs can be limited to a sub unit of the detector readout

  • LI configuration improvements (Lisa)

    • LI runs spanning many runs

    • Tuning the light for individual LEDs

  • RunControl enhancements (Caius, et al)

    • Web status page improvements

    • Fix memory leak (the Sunday problem)


On the horizon

On the Horizon

  • ROP reset and abort implemented through RC

  • The ROP will also calculate sparsification thresholds on pedestal runs (3 widths above pedestal)

  • RC control using new ROP exposed bits

    • Making a diagnosis run sequence can then follow

  • New RC GUI with added feature

    • Upgraded web status page


Fd hv

FD HV

  • Hardware

    • Replaced the faulty ferrups

      • Only three drop outs since January

    • Replaced a HV card

  • Software

    • Ramp changed to hardware ramp

      • No plot yet

    • Changes to warning/reset rules

    • Testing MySQL logging database

      • See DCS section


Fd dcs

FD DCS

  • RPS

    • Replaced the last “old” units

    • Upgrades so fan pack monitoring works (7 of 16 done)

    • Adding smoke detectors for under the shield (per agreement with code officials)

    • Change in logic for loss of power at VFBs (error->warning)

      • Lost a fuse the next week… is it a good idea?

    • Now have a command-line daemon to query the RPS units

  • Enviro

    • Need to get north cluster running & prep for SM2 coil

    • Radon measurements differ wrt CDMS 2 at factor of two

    • Need to replace the north radon counter

  • Coil

    • First field drop out happened last week


State of dcs data

State of DCS Data

  • Erik Beall and rest of DCS working to get the DCS data into it’s own data stream

    • Starting with Nat and the HV

    • Getting data into a local DCS DB

    • Exported into rotorooter / data stream

    • Working on the next step with Robert

    • Then expand across other subsystems

  • Follow on projects

    • Web interface for shifters

      • Current have to do a lot of work to check the DCS in a remote shift

        • Kerberos/ssh for HV, RPS

        • VNC for coil & environmental

        • DAQ tells the grossest problems

    • Making this data available in the offline (core group)

      • Some issues; see computing talk


Fd li system tune up

FD LI – system tune up

  • Test of a UV pulser box

    • Looks better

      • See calibration summary

    • Light level/focus issues

  • LI light level tuning (Hartnell/Falk)

    • LIMaster modified to allow different settings for each LED

    • Jeff has tuned the levels

    • Need to have a better algorithm for spacing the steps


Fee echoes li

FEE Echoes & LI

  • Can’t run like planned now

  • Need ~2 hours for one stability point without data collection

  • See Roy’s talk on the progress

  • Would need to rethink calibration mode if not fixable


Li as a diagnostic tool

LI as a Diagnostic Tool

  • Starting to use LI as a diagnosis tool is was always meant to be

< BOX/LED REPORT >

Missing Pulser Box 14 Completely.

Missing Pulser Box 15 Completely.

Very Quiet Box 12 Only 14464361 Associated Hits (59%)

Very Quiet Box 13 Only 14434576 Associated Hits (59%)

< FEE REPORT >

No Hits Beyond Plane 417

Missing Crate --- Crate 14

Missing Crate --- Crate 15

Very Low Channel/Pixel ---- 0 2 2 0 1 16 -- Plane 55 Only 155 Hits ( 0%)

Very Low Channel/Pixel ---- 2 2 3 1 0 12 -- Plane 122 Only 3117 Hits ( 19%)

Very Low Channel/Pixel ---- 11 1 3 1 2 5 -- Plane 347 Only 0 Hits ( 0%)

Low Chip ---- 12 1 1 1 2 -- Planes 403-405 Only 234789 Hits ( 91%)

Very Low Channel/Pixel ---- 13 2 1 1 0 2 -- Plane 383 Only 207 Hits ( 1%)

Very Low Channel/Pixel ---- 13 2 3 1 1 13 -- Plane 393 Only 40 Hits ( 0%)


Some status plots after recent detector sweep

Some Status Plots (After Recent Detector Sweep)


A more interesting plot the week before

A More “Interesting” Plot(the week before)


Still some more remedial work

Still some more remedial work

  • A number of PIN problems

  • 5 single pixels

  • Few dead LI fibers

Dead Fibre East: 85box 3led 11 Num Wrong: Near 14 Far: 8

Something wrong with East: 226-228box 6led 4 Num Wrong: Near 10 Far: 12

Something wrong with East: 266-268box 8led 18 Num Wrong: Near 12 Far: 16

Something wrong with West: 314-316box 11led 9 Num Wrong: Near 12 Far: 14

Something wrong with East: 339-341box 10led 16 Num Wrong: Near 4 Far: 8

Something wrong with East: 339-341box 10led 20 Num Wrong: Near 10 Far: 6

Something wrong with East: 342-344box 10led 10 Num Wrong: Near 4 Far: 12

Miswired Fibre East: 351box 11led 19 Num Wrong: Near 20 Far: 20

Miswired Fibre East: 351box 11led 20 Num Wrong: Near 16 Far: 18

Dead Fibre East: 363box 10led 7 Num Wrong: Near 14 Far: 8

Dead Fibre East: 375box 11led 20 Num Wrong: Near 16 Far: 16

Miswired Fibre East: 382box 12led 2 Num Wrong: Near 10 Far: 12

Miswired Fibre East: 382box 12led 3 Num Wrong: Near 10 Far: 10

Miswired Fibre East: 382box 12led 5 Num Wrong: Near 10 Far: 10

Miswired Fibre East: 382box 12led 20 Num Wrong: Near 12 Far: 10

Something wrong with East: 402-404box 12led 10 Num Wrong: Near 4 Far: 14

Miswired Fibre East: 407box 13led 7 Num Wrong: Near 20 Far: 20

Miswired Fibre East: 407box 13led 8 Num Wrong: Near 20 Far: 20


Noise problems

Noise Problems

  • We are seeing a recurring but spreading noise problem

  • The affected tubes are essentially dead

  • Problem is the HV connector on the base needs to be reseated

  • Recurring but new ones coming up too

    • Have made 3 passes looking for problems this year

    • ~ half dozen every two weeks

    • 3 have been back each time

  • Needs a fix to make sure these are not continuing to make trouble

    • Doug assigned to the IU MUX guys


Update on electronics mortality dec feb

Update on Electronics Mortality(Dec-Feb)

Dec Jan Feb Tot

VFB swaps 3++ 8+ 0 11

Base swaps 0 9** 5* 14

Cable swaps 1 1 0 2

(1 power, 1 control)

+ 2 cases swapping didn't solve the problem

++ 1 case swapping didn't solve the problem

* All old bases have low R21 (~1kohm)

** 2 cases the swapping didn't fix the problem &2 cases with errors while changing the base


Failure rates

Failure Rates

  • At December Karol told us about the base problem being only being partially to blame for holes (~33% of bases pulled at FarDet and CalDet really bad in UT test stand)

    • Since then Mike led an effort to come up with a revised set of diagnostic steps

    • Results are lower rate of false swaps

    • Base test stand for Soudan later this month

  • Recently Alfons told us similar rates were seen for the VFBs (8 of 23)

    • A test stand for VFBs will be revived at Soudan to reduce the number of false shipments

  • Plans for future

    • FNAL PREP services all front end electronics

    • Test off detector at Soudan before shipping


Numbers spares

Numbers & Spares

  • We evaluated the spares with Roy & Alfons to ensure we have enough to run for many years of operations

  • Total spares seem reasonable … to me

  • Table slightly out of date (Ca. early Feb)


Fd singles rates per plane

FD SinglesRates per plane


Timing system

Timing System

  • There are two features that where originally advertised but are currently not implemented:

    • A precision oscillator in the TCU that would maintain accurate absolute timing if the GPS clock fails

    • The option to generate execute signals with an arbitrary phase with respect to the 40 MHz clock distributed

  • Alfons would like to know if anyone finds the useful enough to pursue

  • Had 1 bad TRC & 1 bad fiber in last 2 months


Magnetics

Magnetics

  • Coil operates routinely

    • Had a supply trip last week for the first time since August

  • Alex Sousa making good progress on calibration control program

    • Still some details to work on

  • Cabling nearly up to date

  • Starting to plan for SM2 coil


Safety at soudan

Safety at Soudan

  • We have had 1safety report at Soudan since (Dec-Feb)

    • Recordable injury related to a previous injury

    • Need to be extra careful with the new crew

    • Reviewing & updating procedures in advance of a refresher course later this month

    • Starting to cross train crew as the number of old timers ramps down

    • Have distributed the new crew among veterans in the crews


Physicists installation

Physicists & Installation

Monthly averages

  • A running total of 999 physicist weeks or 20 physicist years at Soudan on installation tasks

  • Getting there but not done

    • We expect to be commissioning well into the summer

    • There is also a period of training the operations crew later this summer

Today


Needs at soudan

Needs at Soudan

  • We need people (students, postdocs) at either full or 50% time to come up to speed as experts

    • Needed starting in mid to late May

    • Very good summer experience for students who are taking classes or just starting

    • Needs for shift leaders or postdocs to help lead

    • More senior visitors are welcome but already a few so we need to coordinate the dates


Open house

Open House

  • Soudan Open House scheduled for Saturday May, 2nd (the week before fishing opening)

    • Last year we had 650 visitors

    • Need about 10 tour guides

      • Sign up with Bill this weekend

    • More on outreach and summer opportunities/needs from Earl


Operations plan starting this summer

Operations PlanStarting this Summer

  • We are going to test operating for max live time (e.g. like we plan when in beam operations)

    Goal is >>95%

  • Our plan is to staff during 5 days shifts underground

  • We will have two crew (+ hoistman) on call for underground access 24/7 in case of situations that are not remotely rectifiable

    • Will be used by CDMS and MINOS

  • We will operate with local shifts when underground and remote shifts (FNAL + UK?) for other hours


Fd documentation

FD Documentation

http://www-numi.fnal.gov/cgi-bin/noteFDDocsPrint

  • There are currently 26 documents submitted and tagged for the FD manual

    Good guysLaggards

    Front endsEnvironmental Magnetic calibration

    Online monitoring Call lists Safety

    Plane assembly Survey DAQ

    PMT handling Magnet RPS

    Light injection HV

  • Need these for training the operations crew this spring/summer

  • Please increase the priority on this activity to ensure efficient operations this summer/fall

    • Please add in more detailed user’s (idiot’s) guides


Detector operations

Detector operations

  • Since the summer we have been collecting data for calibration and cosmic rays

  • We have also been working to increase the live time

  • Addressing potential hardware impediments

    • HV monitoring & control improved; hardware swaps

    • Optimizing automatic recovery hardware & software

  • Will continue to have 3-4 days per month for DAQ/RPS

  • Hope to get back to some days without taking the detector offline


Exposure summary

Exposure Summary

  • Total exposure (since 1/31/02) 1.52 kty

  • Total fiducial exposure (“ “) 1.04 kty

  • Total fiducial ‘good’ exposure (“ “) 0.94 kty

  • SM1 with field (8/8/02) 0.76 kty

  • Old sec 1 (8/16-11/12/02) 0.18 kty

  • Sec 2 (11/27/02) 0.17 kty

  • Reconfig’d sec 1 (12/?/02) 0.13 kty

  • Lower shield walls (1/27/03)0.09 kty


Cdms 2 update part i from dan bauer

CDMS 2 Update (Part I) (from Dan Bauer)

First tower of detectors installed in “icebox” at Soudan

Second tower installed this week

Radon

purge

Tower 1


Cdms 2 update part ii from dan bauer

CDMS 2 Update (Part II) (from Dan Bauer)

  • Dilution refrigerator and icebox were cooled to 20 mK in Dec, 2002

  • Shielding, DAQ, and Electronics nearly finished


Cdms 2 update part iii from dan bauer

CDMS 2 Update (Part III) (from Dan Bauer)

  • Deep site

    • Much lower neutron background

    • Rapid improvement in sensitivity

  • Expect factor of 10 improvement over CDMS I results by end of 2003

  • Deploy remainder of detectors in 2004 and run until end of 2005 (or longer)

  • Detector performance improvements mean we should still reach original CDMS II goals


Zippy big picture summary

Zippy Big Picture Summary

  • Now 7/8ths complete

  • Will finish this summer

    • Need roughly 8 additional physicist-months late spring and summer

  • Will shift into operations this summer

    • Advanced model for operations

    • Physicists at Soudan beyond summer is an issue


  • Login